OpenStudy (anonymous):

find the critical points of y=x^2(x-2)^2

OpenStudy (anonymous):

may want to start with \[y=x^4-4 x^3+4 x^2\]

OpenStudy (anonymous):

so \[y'=4x^3-12x^2+8x= 4x(x^2-3x+2)\]\]

OpenStudy (anonymous):

then set \[4x(x^2-3x+2)=0\] \[4x(x-2)(x-1)=0\] etc

OpenStudy (anonymous):

thanx satellite73 :-)

OpenStudy (anonymous):

yw